GEn meeting to be held weekly!
Minutes for the GEn weekly meeting on 6th July 2005 - ARC 428(?)
Attendees: Bogdan Wojtsekhowski, Bodo Reitz, Alan Gavalya, Tim Holmstrom, Ameya Kolarkar, Aidan Kelleher, Doug Applegate, Brandon Craver, Albert Shahinyan
~Bodo : Milestones discussion
- Target: Two acceptable cells -> Milestone met.
Anna and Barbara, with ~ 39% and 42% polarizations.
- Neutron Arm: Neutron bars and veto cassettes installed but not yet wired -> 75% of the milestone met.
- BigBite Arm: Calorimeter mounted and tested; wire chambers in Hall tested -> milestone met;
DAQ and veto software -> milestone missed
- Software : Basic Neutron Arm (NA) classes ready but with very rough calibration tools; Olivier working on the pre-shower and shower classes; wire-chamber class ready; basic Analyzer ready.
Wiring on NA and DAQ for the shower need to be done to "meet" the July 1 milestone.
Haven't heard from Syracuse yet. Neil has a task list for the neutron detector, did some of the plots and analysis as suggested by Bogdan. Should forward the results for thresholds for the PMTs to Ameya today. Olivier to manage the shower detector for GEn (he'll transfer the HRS shower class into the GEn package). Ole to look into the BigBite apparatus. Haven't heard from Xiaodong about the BigBite simulations.
~ Brandon : First wire chamber voltage, threshold set up but not yet turned ON. Turning ON the High Voltage (HV) is the plan for today. To start the second chamber soon.
~ Why is the BigBite at 70 degrees? Why not at 50 degrees? Do we need a run-plan for the upcoming test run?
~ Albert : NA assembly
- UVa cassettes installed and housed in cage; 8 cassettes ready for installation; need to prepare 2 CMU cassettes; next - Glasgow detectors and BigBite cabling (high priority).
- Long (100m) delay cables fully ready. About 70% can be connected, top 30% to be done by trained technicians (or train students to do it) since it's too high.
- Glasgow milestone in September.
- Jakob to focus on BB - cable + DAQ.
- 500 cables labeled
- Timing plane + pre-shower to be installed in BB.
~ Tim : New target lab status
- Showed plots from UVa of the spin-up and polarizations measurements for the two GEn cells Anna and Barbara.
- Put in requisition for 2 power supplies for the GEn magnet box.
- Target system should be ready for polarizing a cell till Aug 1; new oven is the main issue.
- To discuss Sasha's extension at JLab with Zein-Eddine.
- To check with Ed Folts about the feasibility of transfer of laser fibers to the Hall in September.
- Ameya to finish setting up optics in new lab by 15th July. JP to help with the optics alignment while Ameya gone.
- Huan working on the NMR system in the old lab while Ameya gone.
- Aidan to present details of the target progress in the next meeting.
~ Alan :
- Mirror assy and detector assy to be ordered this afternoon.
- Mounts completed for the planes (Susan)
- Drive and controller for target under construction (Joyce)
- Target ladder design on hold for a while.
- Oven parts arrive on Jul 27; shop making gluing fixtures.
- Alan going on vacation the end of July.
~ Nilanga :
- Opened chamber 3; several wires noisy -> fixing today.
~ Doug Applegate - New undergrad student from CMU - to work on detector hardware/software
~ John LeRose : Damage control
- Identify problem
- Solve problem
- Separate log entry for things that go wrong (instead of mentioning it only in the shift summary)
- Have fixed keywords for the log entries and spell those out to the shift-takers.
- Have search options on the halog?